# Agentar-Scale-SQL: Advancing Text-to-SQL through Orchestrated Test-Time Scaling

### v4

@@ 1 Introduction
  • Extensive Experiments. Extensive experiments confirm the SOTA performance of Agentar-Scale-SQL. Specifically, Agentar-Scale-SQL achieves an EX of 74.90% on the BIRD development set and 81.67% on the test set, along with an R-VES of 77.00%. With these results [^1] See the official BIRD leaderboard at https://bird-bench.github.io/. The rankings are dated September 28, 2025., Agentar-Scale-SQL ranks first on the BIRD leaderboard. Our findings indicate that scaling is a critical factor for achieving SOTA performance in Text-to-SQL.
@@ 3.3.1 Intrinsic Reasoning SQL Generator
Overview of RL Pipeline. We adopt GRPO (Shao et al., 2024) due to its proven efficiency and effectiveness on structured reasoning tasks. We conduct reinforcement learning on the BIRD training set. Formally, let $\pi_{\theta}$ denote our policy model parameterized by $\theta$. For any given input question $Q_{u}$, its associated evidence and schema, the model generates a set of $N$ candidate SQL queries (i.e., rollouts), $\{o_{Q,1},\dots,o_{Q,N}\}$. Each candidate is then evaluated to yield an explicit reward signal, as described later. This per-input batch of rollouts allows for computing relative advantages, thereby stabilizing learning and fostering robust policy updates.
@@ 4.1 Experimental Setup
Baselines. We compared several top-ranking baseline methods from the overall leaderboard and the single-model leaderboard. The former consists of fifteen baselines, including AskData + GPT-4o (Shkapenyuk et al., 2025), LongData-SQL, CHASE-SQL + Gemini (Pourreza et al., 2025a), JoyDataAgent-SQL, XiYan-SQL (Liu et al., 2025c), among others. The latter comprises nine leading methods, such as Databricks RLVR 32B, Sophon-Text2SQL-32B, Arctic-Text2SQL-R1-32B (Yao et al., 2025).
Metrics. Following prior work (Pourreza et al., 2025a), we use Execution Accuracy (EX), the official metric for the respective leaderboard, as the primary evaluation metric to compare methods. Besides, we adopt the Reward-based Valid Efficiency Score (R-VES) to evaluate the efficiency of the generated SQL.
@@ 4.2 Main Results (RQ1)
As presented in Table 1, our Agentar-Scale-SQL framework establishes a new state-of-the-art (SOTA) on the BIRD benchmark, achieving 81.67% execution accuracy (EX) and 77.00% R-VES on the test set.
This result surpasses the prior SOTA (AskData + GPT-4o) by 0.79% in test EX. Notably, the performance gain is even more pronounced when compared to single-trained models; Agentar-Scale-SQL outperforms the strongest single model (Databricks RLVR 32B) by a substantial 5.99% in test EX. These results empirically validate the effectiveness of our orchestrated scaling strategy.
@@ 5 Conclusion
Despite its effectiveness, Agentar-Scale-SQL’s reliance on orchestrated test-time scaling introduces several key limitations. The framework’s primary drawback is its substantial computational overhead and high latency due to the multiple LLM calls for generation, refinement, and selection, making it less suitable for real-time applications. Furthermore, its performance is fundamentally bounded by the capabilities of the underlying base LLM, and it is susceptible to cascading errors where a failure in an early stage, such as task understanding, can compromise the entire process.
@@ Appendix A BIRD Leaderboard: Bar Chart and Snapshot
We present the BIRD rankings as a bar chart (Figure 7) and a snapshot (Figure 8) as of September 28, 2025.
@@
Figure 7: BIRD leaderboard: EX and R-VES performance comparison.
Figure 8: A snapshot of the BIRD benchmark’s dynamic leaderboard as of September 28, 2025.
@@ Appendix D Discussion of Schema Linking
Appendix E Prompts
@@ 4. You need to enclose the final SQL query in a code block using the following format:
Appendix F Further Work

### v5

@@ 1 Introduction
  • Extensive Experiments. Extensive experiments confirm the SOTA performance of Agentar-Scale-SQL. Specifically, Agentar-Scale-SQL achieves an EX of 74.90% on the BIRD development set and 81.67% on the test set, along with an R-VES of 77.00%. With these results [^1] See the official BIRD leaderboard at https://bird-bench.github.io/. The rankings are dated November 27, 2025., Agentar-Scale-SQL ranks first on the BIRD leaderboard. Our findings indicate that scaling is a critical factor for achieving SOTA performance in Text-to-SQL.
@@ 3.3.1 Intrinsic Reasoning SQL Generator
Overview of RL Pipeline. We adopt GRPO (Shao et al., 2024) due to its proven efficiency and effectiveness on structured reasoning tasks. Unlike Arctic-Text2SQL-R1, we exclusively use the training data from the BIRD dataset, without any extra data. Formally, let $\pi_{\theta}$ denote our policy model parameterized by $\theta$. For any given input question $Q_{u}$, its associated evidence and schema, the model generates a set of $N$ candidate SQL queries (i.e., rollouts), $\{o_{Q,1},\dots,o_{Q,N}\}$. Each candidate is then evaluated to yield an explicit reward signal, as described later. This per-input batch of rollouts allows for computing relative advantages, thereby stabilizing learning and fostering robust policy updates.
@@ 4.1 Experimental Setup
Baselines. We compared several top-ranking baseline methods from the overall leaderboard and the single-model leaderboard. The former consists of fifteen baselines, including AskData + GPT-4o (Shkapenyuk et al., 2025), LongData-SQL, CHASE-SQL + Gemini (Pourreza et al., 2025a), JoyDataAgent-SQL, XiYan-SQL (Liu et al., 2025c), among others. The latter comprises eight leading methods, such as Gemini-SQL, Databricks RLVR 32B, Sophon-Text2SQL-32B, Arctic-Text2SQL-R1-32B (Yao et al., 2025).
Metrics. Following prior work (Pourreza et al., 2025a), we use Execution Accuracy (EX), the official metric for the respective leaderboard, as the primary evaluation metric to compare methods. Besides, we adopt the official Reward-based Valid Efficiency Score (R-VES) to evaluate the efficiency of the generated SQL.
@@ 4.2 Main Results (RQ1)
As presented in Table 1, our Agentar-Scale-SQL framework establishes a new SOTA on the BIRD benchmark, achieving 81.67% execution accuracy (EX) and 77.00% R-VES on the test set. This result surpasses the prior SOTA (AskData + GPT-4o) by 0.79% in test EX. It is worth noting that these approaches employ powerful models, such as Gemini 2.5 Pro and Claude 4.
Notably, the performance gain is even more pronounced when compared to single-trained models; Agentar-Scale-SQL outperforms the strongest single model (Gemini-SQL) by a substantial 5.54% in test EX. These results empirically validate the effectiveness of our orchestrated scaling strategy.
@@ 5 Conclusion
Despite its effectiveness, Agentar-Scale-SQL’s reliance on orchestrated test-time scaling introduces several key limitations. The framework’s primary drawback is its substantial computational overhead and high latency due to the multiple LLM calls for generation, refinement, and selection, making it less suitable for real-time applications. In our commercial practice (developing B2B ChatBI products), we observe that enterprise clients prioritize Accuracy above all else. In decision-making scenarios, a hallucinatory SQL query is unacceptable, whereas a latency of several seconds is often tolerable for complex analytics generation. Agentar-Scale-SQL is designed specifically to bridge the "last mile" of accuracy for these production requirements. Furthermore, its performance is fundamentally bounded by the capabilities of the underlying base LLM, and it is susceptible to cascading errors where a failure in an early stage, such as task understanding, can compromise the entire process.
@@ Appendix A BIRD Leaderboard: Bar Chart and Snapshot
We present the BIRD rankings as a bar chart (Figure 7) and a snapshot (Figure 8).
@@
Figure 7: BIRD leaderboard (as of September 28, 2025): EX and R-VES performance comparison.
Figure 8: A snapshot of the BIRD benchmark’s dynamic leaderboard as of November 27, 2025.
@@ Appendix D Discussion of Schema Linking
Appendix E Extended Analysis: Scaling, Efficiency, and Comparisons
In this section, we provide further elaboration on the core contributions of our method regarding test-time scaling orchestration, a detailed comparison of scaling capabilities against SOTA baselines, and a quantitative analysis of computational costs versus utility.
E.1 Orchestrated Test-Time Scaling
Our work distinguishes itself from naive combinations of existing modules through a systematic orchestration of Test-Time Scaling across three distinct dimensions. Unlike simple stacking, we introduce specific designs to maximize synergy, such as Diverse Synthesis (combining breadth via ICL with depth via RL-tuned reasoning) and Tournament Selection (an RL-enhanced pairwise judge). This orchestration allows Agentar-Scale-SQL to validate the “Scaling Hypothesis” in Text-to-SQL, outperforming the strongest single model (Gemini-SQL) by 5.54% and the previous SOTA (AskData + GPT-4o) by 0.79%.
E.2 Comparison of Scaling Capabilities
Table 4 categorizes existing SOTA methods based on their supported scaling dimensions. While most frameworks support only one or two dimensions, our method is the first to fully integrate all three, offering a comprehensive optimization space.
Table 4: Comparison of scaling capabilities across SOTA methods. Our method uniquely supports Internal, Sequential, and Parallel scaling simultaneously.
Methods Internal Scaling Sequential Scaling Parallel Scaling
Overall
XiYan-SQL -
Contextual-SQL - -
CHASE-SQL + Gemini -
AskData + GPT-4o - -
Single Trained Models
Arctic-Text2SQL-R1-32B - -
Databricks RLVR 32B - -
Gemini-SQL - -
Agentar-Scale-SQL (Ours)
E.3 Computational Cost vs. Practical Utility
Table 5: Comparison of inference costs (estimated number of LLM calls).
Methods Task Understanding Generation Refinement Selection Total Calls
Overall
Contextual-SQL - 1024 - 1024 2048
CHASE-SQL + Gemini 32 16 32 10 297
Single Trained Models
Sophon-Text2SQL-32B - 8 - 32 - - 8 - 32
Databricks RLVR 32B - 1 - 7 - - 1 - 7
Gemini-SQL - 1 - 7 - - 1 - 7
Agentar-Scale-SQL (Ours) 1 9 45 - 153 1 - 9 56 - 172
While test-time scaling inherently introduces higher latency compared to single-pass models, this is a deliberate trade-off designed for high-stakes enterprise scenarios. In B2B ChatBI products, accuracy is paramount; a hallucinated query in a financial report is unacceptable, whereas a latency of 10–30 seconds is often tolerable for complex analytics.
Furthermore, our framework offers high configurability. Users can dynamically adjust the number of candidates ($N$) or skip the refinement stage to reduce costs for simpler queries, bridging the “last mile” of accuracy only when necessary.
Table 5 provides an approximate comparison of inference costs using the number of LLM calls as a metric. Although our method involves more steps than single-pass models, it significantly optimizes the total calls compared to other multi-stage frameworks (e.g., Contextual-SQL) while achieving SOTA performance.
Appendix F Prompts
@@ 4. You need to enclose the final SQL query in a code block using the following format:
Appendix G Further Work